### v1

# Emergent synchrony in oscillator networks with adaptive arbitrary-order interactions

Abstract
Dynamics of complex systems are often driven by interactions that extend beyond pairwise links, underscoring the need to establish a correspondence between interpretable system parameters and emergent phenomena in hypergraph-based networks. The current work formulates an adaptive Kuramoto model that incorporates hyperedges of arbitrary order and explores their effects on synchronization. By deriving the exact order parameter dynamics in the thermodynamic limit, analytical expressions governing the collective dynamics are obtained. Subsequent numerics confirm the analytical predictions, in addition to capturing qualitatively different dynamical regimes and phase transitions. Further investigations based on order parameter distributions demonstrate how fluctuations, arising due to finite system size, can influence the long-term system dynamics. These results provide important insights and can have diverse applications, such as designing optimal surgical procedures for drug-resistant epilepsy and identifying the sources of rumours in a social network.
preprint: APS/123-QED
Introduction — Recent studies across different disciplines show that complex systems involve links that extend beyond pairwise edges and include hyperedges that capture interactions between irreducible groups of three or more nodes simultaneously [hypergraphtheoryBretto2013, majhi2022dynamics, shi2022simplicial, boccaletti2023structure, PhysRevLett.131.207401]; see Fig. 1(a) for a sample schematic. Interactions in social dynamics are an archetypal example and include collaborations in scientific research [lungeanu2021team, moore2012analyzing], group interactions in opinion spreading [wang2020social, ghahremani2023novel], and dynamics of consensus [PhysRevE.101.032310, PhysRevE.104.064305], among others. In biological systems, which are often replete with higher-order interactions, neuroscience stands out as a promising domain [giusti2016two, hindriks2025unraveling]. Higher-order connectivity is evident in the human brain [karmelic2022emergent] (also see Fig. 1(b)), and has been linked to avalanches and other neural dynamics [yu2011higher, millan2025spatio, biswas2025effect], and could provide deeper insights into pathological states [PhysRevResearch.5.013074]. These also appear in models describing chemical reactions [wen2023chemical, brauch2013higher], power grids [ghasemi2022higher, PhysRevE.109.024212], and the evolution of ecological systems [swain2022higher, gallardo2024higher, bairey2016high, PhysRevE.111.014309], highlighting their ubiquity across domains.
Figure 1: (a): Schematic representation of hyperedges in the brain: In (i), pink circles indicate nodes, or Broadman areas of functional significance, whereas the lines (edges) and coloured regions (hyperedges) denote pairwise and higher-order interactions, respectively. These can appear at multiple spatial scales, such as macroscopic functional connectivity patterns from fMRI-BOLD time series and microscopic neuronal scale, shown in (ii) and (iii), respectively; (b): Number of interactions in a brain network, as a function of its order, relative to pairwise interactions [karmelic2022emergent, cabral2022metastable]; (c): Temporal variation of intracranial (iEEG) signals from different channels [bougou2025mesoscale, dataset], with the histogram (top-inset) and time series (bottom-inset) of coherence, highlighting its increase during epilepsy, the onset of which is marked with a vertical red line.
The dynamics of systems with higher-order interactions can be described by equations of the form
$\frac{{\rm d}x_{i}}{{\rm d}t}$ $=f_{i}(x_{i})+\sum_{\mathclap{d=1;j_{1},..,j_{d}=1}}^{D;N}A^{(d)}_{i,j_{1},..,j_{d}}g^{(d)}(x_{i},x_{j_{1}},..,x_{j_{d}}),$ (1)
where $N$ and $D$ denote the number of constituent nodes and the maximum order of interaction, respectively. Oscillations are a characteristic feature often observed in such systems [rabinovich2012oscillations], with their synchronization representing a vital emergent phenomenon having important implications across disciplines [arenas2008synchronization]; for example, epilepsy is often characterized by synchronous activity in the brain [osterhage2007measuring, lehnertz2009synchronization]; see Fig. 1(c) for details. To this end, the Kuramoto model of coupled phase oscillators [RevModPhys.77.137, gupta2014kuramoto] serves as a paradigmatic framework for quantitatively exploring synchronization in a systemic manner. To capture realistic systems, it has also been extensively modified with features such as complex connectivity matrices [jain2023composed, chen2023critical, peron2019onset, tong2018exponential], symmetry-breaking mechanisms [manoranjani2023generalization, biswas2024symmetry], time-delayed interactions [yeung1999time, wu2018dynamics, ross2021dynamics, budzinski2023analytical], higher-order couplings [xu2021spectrum, costa2024bifurcations, costa2025exact, moyal2024rotating, rajwani2023tiered, carballosa2023cluster, sabhahit2024prolonged, skardal2020higher], as well as different combinations of each of these [skardal2022tiered]. These find applications in a myriad of different domains [guo2021overviews, wang2021dynamic, blanter2016kuramoto, fioriti2008stability], and are of particular importance in neuroscience, where they provide a tractable mathematical framework to model complex neuro-oscillatory phenomena [odor2019critical, heggli2019kuramoto, PhysRevLett.120.244101, odor2021effect, pathak2022biophysical, ranjan2024propagation, PhysRevE.111.044310].
This work investigates an adaptive Kuramoto model, which incorporates arbitrary-order interactions of both symmetric and asymmetric types, to explore the effects of higher-order interactions and their interplay with synchronization. The dynamics of the system are assumed to be governed by equations of the form
$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r_{1})\epsilon_{1}}{N}\sum_{j_{1}=1}^{N}\sin\Big(\phi_{j_{1}}-\phi_{i}-\delta_{1}\Big)+\\ \sum_{d=2}^{d=D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{N^{d}}\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(d\phi_{j_{d}}-\sum_{k=1}^{d-1}\phi_{j_{k}}-\phi_{i}-\delta_{d}\right)+\right.\\ \left.\frac{f_{d}^{\prime}(r_{1})\epsilon^{\prime}_{d}}{N^{d}}\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(d\phi_{j_{d}}+\sum_{k=1}^{d-1}\phi_{j_{k}}-\phi_{i}-\delta^{\prime}_{d}\right)\right],$ (2)
where $z_{k}=r_{k}e^{{\rm i}\psi_{k}}=N^{-1}\sum_{j=1}^{N}e^{{\rm i}k\phi_{j}}\ \forall\ k\geq 1$ are the Kuramoto-Daido order parameters [clusella2020irregular] and ${\rm i}=\sqrt{-1}$. In this model, the interaction terms are strongly influenced by the corresponding adaptation functions, the inclusion of which represents a fundamental dynamical feature of natural systems [sawicki2023perspectives]. Adaptation allows the characteristic properties of the system to change in response to the collective state of the system and is seen across bio-physical, socio-economic, and engineering systems, and is particularly prevalent in the study of neural activity; for example, spike-frequency adaptation models the reduction in neural firing arising from the fatigue of continuous spiking activity. [PhysRevE.107.024311]. In the context of synchronization, this allows the modulation of the global coupling strength in response to the extent of synchronization in the system [jin2023synchronization], thereby allowing the system to reach a desired state of synchronicity. A universal and well-investigated form of the adaptation function is a power-law, given by $f(r_{1})\propto r_{1}^{\alpha}$, where $\alpha$ controls the nature of the feedback and the resulting dynamics [PhysRevE.75.017201, PhysRevE.102.012219, biswas2024effect].
Analysis — For an infinitely large number of oscillators, an assumption relaxed later, the state of the system can be described by a probability density function $F(\phi,\omega;t)$, which satisfies a continuity equation of the form $\partial_{t}F+\partial_{\phi}(F\dot{\phi})=0$. Since $F(\phi+2\pi,\omega;t)=F(\phi,\omega;t)$, it can be expanded in a Fourier series in $\phi$, where $F_{n}(\omega,t)$ is the $n^{\rm th}$ Fourier coefficient and $F_{-n}(\omega,t)=F^{*}_{n}(\omega,t)$. Using the Ott-Antonsen ansatz [ott2008low, ott2009long], it can be shown that
$\frac{\partial v}{\partial t}+{\rm i}\omega v+\frac{f_{1}(r_{1})\epsilon_{1}}{2}\left(z_{1}e^{-{\rm i}\delta_{1}}v^{2}-z_{1}^{*}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}\delta_{d}}v^{2}-z_{d}^{*}z_{1}^{d-1}e^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}\delta^{\prime}_{d}}v^{2}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}\delta^{\prime}_{d}}\right)\right]=0,$ (3)
where $F_{n}(\omega,t)=v^{n}(\omega,t)$. Subsequently, by assuming that $\omega_{i}$ are sampled from a Cauchy distribution with mean $\omega_{0}$ and scale factor $\Delta$, the order parameters are obtained as $z_{d}=\left[v^{*}(\omega_{0}-i\Delta,t)\right]^{d}$. Substituting into Eq. (3) and re-labelling $z_{1}$ as $z$ for simplicity, it takes the form
$\frac{\partial z}{\partial t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{f_{1}(r)\epsilon_{1}}{2}\left(ze^{-{\rm i}\delta_{1}}-z|z|^{2}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r)\epsilon_{d}}{2}\left(|z|^{2(d-1)}ze^{-{\rm i}\delta_{d}}-|z|^{2d}ze^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r)\epsilon^{\prime}_{d}}{2}\left(z^{2d-1}e^{-{\rm i}\delta^{\prime}_{d}}-|z|^{4}(z^{*})^{2d-3}e^{{\rm i}\delta^{\prime}_{d}}\right)\right].$ (4)
Evidently, the incoherent state is stable only for $\epsilon_{1}\leq\epsilon_{1}^{c}=2\Delta/f_{1}(0)\cos(\delta_{1})$, which denotes the transition point from an asynchronous state to a synchronous state. In contrast, the dynamics of the synchronous states are governed by equations of the form
$\frac{{\rm d}r}{{\rm d}t}=R(r,\psi)=-\Delta r+r(1-r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\cos(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\cos(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\cos(\chi^{\prime}_{d})\right)\right],$ (5)
and
$\frac{{\rm d}\psi}{{\rm d}t}=S(r,\psi)=\omega_{0}-(1+r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\sin(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\sin(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\sin(\chi^{\prime}_{d})\right)\right],$ (6)
where $\chi^{\prime}_{d}=\delta^{\prime}_{d}-2(d-1)\psi$. These two equations dictate the dynamics of $r(t)$ and $\psi(t)$, the steady state solution $(r_{s},\psi_{s})$, and the subsequent stability analysis provides insights into the overall system dynamics.
(a) $\epsilon_{1}=0.6$
(b) $\epsilon_{1}=1.3$
(c) $\epsilon_{1}=1.8$
(d) $\epsilon_{1}=2.6$
Figure 2: (a)-(d): Left: Graphical solution of Eq. (10). The black and magenta curves denotes the LHS and RHS, respectively, whereas the green and red broken vertical lines (in inset) denote the stable and unstable steady-state solutions respectively, as determined from the eigenvalues of the Jacobian of Eqs. (8) and (9); Right: Temporal variation of order parameter $r(t)$. The blue and red curves denote the asynchronous and synchronous initial conditions, respectively, whereas the horizontal line denotes $r_{s}$. For all figures, $\omega_{0}=1.5$, $\gamma=0.2$, and $\epsilon_{2}=5.0$.
Results — As a relatively simple example of Eq. (2), a system described by equations of the form
$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r)\epsilon_{1}}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i})+\\ \frac{f_{2}(r)\epsilon_{2}}{N^{2}}\sum_{j,k=1}^{N}\Bigg[\sin(2\phi_{k}-\phi_{j}-\phi_{i})+\xi\sin(2\phi_{k}+\phi_{j}-\phi_{i})\Bigg],$ (7)
is investigated, where $\xi=\epsilon_{2}^{\prime}/\epsilon_{2}$. For the sake of simplicity, $f_{d}(r)$ are assumed to be of the form $f_{d}(r)=(1+\gamma r)^{\pm 2}$ for $d=1$ and $d=2$, respectively, whereas $\gamma$ is a constant. In this case, Eqs. (5) and (6) take the form
$\dot{r}$ $=-\Delta r+r(1-r^{2})\Big[\sigma_{1}(r)+\sigma_{2}(r)\Big(1+\xi\cos(2\psi)\Big)\Big],$ (8)
$\dot{\psi}$ $=\omega_{0}+\xi(1+r^{2})\sigma_{2}(r)\sin(2\psi),$ (9)
respectively, where $\sigma_{n}(r)=f_{n}(r)\epsilon_{n}r^{2(n-1)}/2$ for $n=1,2$, and $\epsilon_{1}^{c}=2\Delta$. The value of $r_{s}$ can be obtained as a solution of the algebraic equation
$\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right]^{2}+\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right]^{2}=\xi^{2}\sigma_{2}^{2}(r_{s}).$ (10)
However, in most cases, it is not possible to find closed-form analytical solutions to the above equations, thus requiring the use of numerical techniques.
For the rest of the work, $(\Delta,\xi)=(1.0,0.4)$ in Eq. (7) for the sake of simplicity. Furthermore, as an example of adaptive systems, the specific case of $\gamma=0.2$ is demonstrated in Fig. 2, for different values of $\epsilon_{1}$, whereas the value of $\omega_{0}$ and $\epsilon_{2}$ is set at $1.5$ and $5.0$, respectively  [^1] The non-adaptive case (i.e., $\gamma=0$) has already been studied in Ref. [biswas2024symmetry].. For $\epsilon_{1}=0.6$, it is evident that Eq. (10) admits no solutions. Furthermore, the asynchronous state is stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system always converges towards $r(t)\approx 0$, irrespective of initial conditions; see Fig. 2(a). However, as the value of $\epsilon_{1}$ is increased to $1.3$, there is a qualitative change in the dynamics. In this case, the asynchronous state remains stable, since $\epsilon_{1}<\epsilon_{1}^{c}$, and is evident from the temporal variation of $r(t)$, which converges towards $r(t)\approx 0$ when initialised in an asynchronous state. However, the system also demonstrates an unsteady synchronised state, where the order parameter shows oscillations about a mean value; see Fig. 2(b).
(a) $\gamma=0.05$
(b) $\gamma=0.10$
(c) $\gamma=0.15$
(d) $\gamma=0.20$
(e) $\gamma=0.25$
(f) $\epsilon_{2}=2.0$
(g) $\epsilon_{2}=3.0$
(h) $\epsilon_{2}=4.0$
(i) $\epsilon_{2}=5.0$
(j) $\epsilon_{2}=6.0$
Figure 3: (a)-(j): Variation of steady state order parameter $r$ as a function of $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ for different values of $\gamma$ (top-row; $\epsilon_{2}=5.0$), and $\epsilon_{2}$ (bottom-row; $\gamma=0.2$). The vertical line corresponds to the forward transition point (i.e., $E^{c}=\epsilon_{1}^{c}/\epsilon_{2}$), whereas the blue and red curves denote the forward and backward variation of $E$, respectively.
As demonstrated in Fig. 2(c), increasing $\epsilon_{1}$ to $1.8$ entails further change in the dynamics. In this case, Eq. (10) admits two solutions $r_{s}=0.817$ and $r_{s}=0.849$. The eigenvalue analysis of these solutions demonstrates that the former is stable, whereas the latter is unstable in nature. However, the asynchronous state is also stable since $\epsilon_{1}<\epsilon_{1}^{c}$. Thus, the system is bistable, where the temporal variation of $r(t)$ reaches two different steady states depending on the initial condition. In particular, the asynchronous initial conditions continue to be asynchronous, whereas the synchronous initial condition converges to $r(t)\approx 0.82$. Upon further increasing $\epsilon_{1}$ to $2.6$, the asynchronous state loses stability since $\epsilon_{1}>\epsilon_{1}^{c}$, as demonstrated in Fig. 2(d). However, Eq. (10) continues to have two solutions, one at $r_{s}=0.821$, which is stable, and the other at $r_{s}=0.872$, which is unstable, as determined by the eigenvalue analysis. Thus, the system is now monostable, which is demonstrated by the temporal variation of $r(t)$, which converges around $r(t)\approx 0.82$ irrespective of the initial state of the system.
Relative effect of interactions: To quantify the relative effect of pairwise and higher-order interactions, the dynamics of the system is investigated via the variation in the steady state value of the order parameter as a function of the ratio $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ in Fig. 3, representing the relative strength of the pairwise and triadic interactions. The top and bottom rows correspond to $\epsilon_{2}=5.0$ and $\gamma=0.2$, respectively. In the former, the value of $\gamma$ varies in the range $[0.05,0.25]$ in steps of $0.05$ across the sub-figures, whereas in the latter, the value of $\epsilon_{2}$ varies in the range $[2.0,6.0]$ in steps of $1.0$.
In Figs. 3(a)-3(e), upon increasing $\gamma$, the transition to synchrony upon forward variation of $E$ is observed to be explosive, which is characterized by a sudden increase in $r$ at the critical point $E\approx 0.4$. In contrast, the backward transition point, captured by the sharp drop in $r$ upon reducing $E$, reduces with increasing values of $\gamma$. However, the post-transition state upon increasing $E$, as well as the route to desynchronization upon reducing $E$, varies for different values of $\gamma$. In particular, for $\gamma=0.05$, the system demonstrates an unsteady synchronous state characterized by an oscillatory temporal variation of $r(t)$ immediately after the explosive transition. Furthermore, tuning $E$ beyond $1.0$ results in a steady synchronous state; see Fig. 3(a). Subsequently, upon increasing $\gamma$ to $0.10$ and $0.15$, the region of $E$ for which the system demonstrates an unsteady synchronous state reduces, resulting in a larger parameter range for which the steady synchronous state is observed; see Figs. 3(b) and 3(c). Furthermore, upon increasing $\gamma$ to $0.20$ and $0.25$, the system exhibits only a steady synchronous state after the transition. However, a key difference is that, in the former, the system exhibits an unsteady synchronous state in a portion of the bi-stable regime before desynchronization, whereas the latter maintains the steady synchronous state enroute to desynchronization; see Figs. 3(d) and 3(e).
The curves in Figs. 3(f)-3(j) demonstrate the qualitative dynamical changes upon changing $\epsilon_{2}$. Unlike the previous case, changing $\epsilon_{2}$ modifies the characteristic nature of the transition to synchrony upon varying $E$. In particular, while for $\epsilon_{2}=2.0$, the transition is observed to be continuos with a smooth increase in the steady-state value of $r$ beyond the critical point (see Fig. 3(f)), higher values of $\epsilon_{2}$ result in explosive synchronization transitions, in addition to the formation of a bi-stable zone; see Figs. 3(g)-3(j). Similar to the previous case, the post-transition synchronous states can be both unsteady or steady (for example, see Figs. 3(h) and 3(i)), whereas, for $\epsilon_{2}=6.0$, the system does not undergo any de-synchronisation transition upon reducing $E$; see Fig. 3(j).
Effect of system size: Real bio-physical systems, such as the brain, are typically modelled as a connected system with a finite number of constituent dynamical units (here, oscillators), the size of which could be of different orders depending on the level of abstraction. Thus, while the previous sections demonstrate steady-state results for an infinitely large number of oscillators, the synchronization phenomena in finite-sized systems are typically affected by fluctuations which can be better characterised by constructing a probability density of the order parameter, denoted by $p(r)$. This is computed by simulating the system for a large number of time-steps and repeated over multiple initial conditions, as well as for two qualitatively different sets characterised by $r(0)\approx 0.0$ (i.e., nearly asynchronous) and $r(0)\approx 1.0$ (i.e., highly synchronous), respectively, with the results demonstrated in Fig. 4 for $N=50$.
(a) $\epsilon_{1}=0.6$
(b) $\epsilon_{1}=1.3$
(c) $\epsilon_{1}=1.8$
(d) $\epsilon_{1}=2.6$
Figure 4: (a)-(d): Distribution of order parameter $p(r)$, collapsed across time, for $N=50$, and different values of $\epsilon_{1}$. The blue and red curves denote completely asynchronous and synchronous initial conditions, respectively, whereas the vertical line denotes $r_{s}$. For all figures, $(\omega_{0},\epsilon_{2},\gamma)=(1.5,5.0,0.2)$.
It is observed that, for $\epsilon_{1}=0.6$ in Fig. 4(a), both distributions show a bimodal structure, with peaks near $r\approx 0.1$ and $r\approx 0.8$ respectively. However, the blue curve is substantially skewed towards $r=0$, whereas the red curve has two peaks of similar height. This indicates that, if initialized asynchronously, there is a higher probability of the system remaining asynchronous compared to transitioning to a highly synchronous state. On the other hand, if the system is initialized in a highly synchronous state, there is a nearly equal probability that the system will transition to an asynchronous state or remain in a synchronous state.
In comparison, for $\epsilon_{1}=1.3$ in Fig. 4(b), both distributions retain their bimodal structure. However, in contrast to the previous case, both distributions are skewed heavily towards $r=1$. Here, the red curve is much more sharply peaked near $r\approx 0.8$ compared to $r\approx 0.1$, indicating a higher probability of finding the system in a highly synchronous state if the system is initialised in a synchronous state. In comparison, the blue curve is only slightly skewed towards the synchronous state, indicating that the probability of finding the system in a synchronous state is marginally higher than that of finding it in an asynchronous state if the system is initialised in an asynchronous state. Finally, for $\epsilon_{1}=1.8$ and $2.6$, the distributions become nearly unimodal and skewed towards $r=1$, with a peak near $r\approx 0.8$. This implies that, for any initial condition, there is a much higher probability of observing the system to be in a highly synchronous state; see Figs. 4(c) and 4(d).
Conclusions — This work investigates the emergence of synchrony in a population of Kuramoto oscillators, where the interactions are both adaptive and higher-order in nature, having both rotational symmetry-preserving and symmetry-breaking components. To this end, exact analytical results for the synchronization states, their stability, and the corresponding phase transition points are obtained in the thermodynamic limit for any order of interaction. Subsequently, these were verified through large-scale simulations, focusing on the effects of triadic interactions, revealing a repertoire of dynamical states, including monostable, bistable, and unstable cases. Furthermore, varying the relative coupling strength between pairwise and higher-order interactions shows a qualitative change in the system dynamics, including a scenario where the desynchronization transition is suppressed. The effects of finite system size were also examined through numerically constructed order parameter distributions, highlighting how fluctuations in the temporal variation of the order parameter, arising out of the finite size of the system, can lead to transitions to a highly synchronous state that would not otherwise occur in the thermodynamic limit.
The results presented provide important insights into the interplay of higher-order interactions and other coupling modalities, allowing for more precise delineation of transition boundaries, yielding a comprehensive “map” of the accessible states. A notable application lies in understanding epileptic seizures, where incorporating higher-order interactions could improve surgical strategies for drug-resistant epilepsy [schramm2008temporal, jirsa2023personalised, wang2023delineating] by revealing underlying epileptogenic foci that can remain obscured in purely pairwise models, thereby facilitating precise and targeted interventions. Another important example concerns the dissemination of information in social systems, where analyzing higher-order connectivity can uncover underlying “cells”, i.e., individuals or organizations that act as synchronized sources of misinformation, and could explain why misleading content often propagates far more rapidly than verified information [vosoughi2018spread]. Future work could extend this framework to involve empirical networks that include hyper-graph connectivities (for example, see Ref. [ko2022growth]), integrate domain-specific models beyond the generic Kuramoto framework adopted here, and estimate parameters using available datasets to obtain specific insights.
Acknowledgements — The authors acknowledge the financial support provided via NBRC Core Funds. DB acknowledges discussions with Dr. Proloy Das (NBRC), PV Vinitha (NBRC), and Shashwati Tripathi (NBRC).
Declaration of competing interest — The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Appendix
Varying system size — In Fig. 5, the figures highlight the characteristic change in the order parameter dynamics as the system size $N$ is increased for different values of the pairwise coupling strength $\epsilon_{1}$.
Figure 5: (a)-(l): $p(r)$ for different values of $(N,\epsilon_{1})$, where $N=100$, $500$, $1000$ (left to right), and $\epsilon_{1}=0.6$, $1.3$, $1.8$, $2.6$ (top to bottom). Here, $(\omega_{0},\epsilon_{2},\gamma)=(1.5,5.0,0.2)$, and the colours carry the same meaning as Fig. 4.
The results demonstrate that, for values of $\epsilon_{1}$ given by $0.6$, $1.3$, and $1.8$, the blue curve (asynchronous initial conditions) loses its bi-modal structure and becomes sharply peaked near $r=0$ as the value of $N$ is increased; see Fig. 5(a)-5(i). In contrast, for $\epsilon_{1}=2.6$, the blue curve becomes sharply peaked near $r\approx 0.8$ upon increasing $N$; see Figs. 5(j)-5(l). These observations are consistent with the fact that, in the limit $N\rightarrow\infty$, the system has a stable fixed point at $r=0$ $\forall$ $\epsilon_{1}\leq 2$. The observations regarding the red curve (synchronous initial conditions) are markedly different. In particular, for $\epsilon_{1}=0.6$, the red curve completely loses its bi-modal structure and becomes nearly identical to the blue curve with increasing $N$. This is consistent with previous analysis, which showed that $r=0$ is the only steady synchronous state for this parameter value in the case $N\rightarrow\infty$. However, upon increasing $N$ for $\epsilon_{1}=1.3$, the red curve becomes locally bi-modal around $r\approx 0.8$. This indicates the presence of oscillatory dynamics in the order parameter and is consistent with the previous findings. Finally, upon increasing $\epsilon_{1}$ to $1.8$ and $2.6$, the red curve becomes unimodal with a sharp peak near $r\approx 0.8$ as the value of $N$ is increased, which is indicative of a stable synchronous fixed point.
Simulation details — A direct simulation of the Kuramoto model with a coupling term of the form
$I=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}+\sigma\phi_{i}+\alpha\right)$ (11)
would be prohibitively time-consuming due to the presence of nested summations, especially for $d\geq 2$. To address this, the RHS of Eq. (11) is expanded as $I=X_{d}\cos(\sigma\phi_{i}+\alpha)+Y_{d}\sin(\sigma\phi_{i}+\alpha)$, where
$X_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right),\ Y_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\cos\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right).$ (12)
Therefore, $X_{d}$ can be expanded as
$X_{d}=\sum_{j_{1},..,j_{d}=1}^{N}\left[\sin\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\cos(c_{d}\phi_{j_{d}})+\right.\\ \left.\cos\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\sin(c_{d}\phi_{j_{d}})\right],$ (13)
which implies $X_{d}=X_{d-1}C_{d}+Y_{d-1}S_{d}$ and
$C_{d}=\sum_{j=1}^{N}\cos(c_{d}\phi_{j}),\ S_{d}=\sum_{j=1}^{N}\sin(c_{d}\phi_{j}).$ (14)
Similarly, $Y_{d}=Y_{d-1}C_{d}-X_{d-1}S_{d}$, whereas $(X_{1},Y_{1})=(S_{1},C_{1})$. This provides an iterative method to efficiently compute the interaction terms of the form of Eq. (11).

### v2

@@
 Abstract
-Dynamics of complex systems are often driven by interactions that extend beyond pairwise links, underscoring the need to establish a correspondence between interpretable system parameters and emergent phenomena in hypergraph-based networks.
-The current work formulates an adaptive Kuramoto model that incorporates hyperedges of arbitrary order and explores their effects on synchronization.
+Complex physical systems are often governed by interactions that extend beyond pairwise links, underscoring the need to establish a map between interpretable system parameters and emergent synchronisation phenomena in hyper-graphs.
+To achieve this, the current work formulates an adaptive Kuramoto model that incorporates hyperedges of arbitrary order and explores their effects on synchronisation.
 By deriving the exact order parameter dynamics in the thermodynamic limit, analytical expressions governing the collective dynamics are obtained.
 Subsequent numerics confirm the analytical predictions, in addition to capturing qualitatively different dynamical regimes and phase transitions.
-Further investigations based on order parameter distributions demonstrate how fluctuations, arising due to finite system size, can influence the long-term system dynamics.
-These results provide important insights and can have diverse applications, such as designing optimal surgical procedures for drug-resistant epilepsy and identifying the sources of rumours in a social network.
-preprint: APS/123-QED
-Introduction — Recent studies across different disciplines show that complex systems involve links that extend beyond pairwise edges and include hyperedges that capture interactions between irreducible groups of three or more nodes simultaneously [hypergraphtheoryBretto2013, majhi2022dynamics, shi2022simplicial, boccaletti2023structure, PhysRevLett.131.207401]; see Fig.
+Further investigations based on numerically constructed order parameter distributions demonstrate how fluctuations due to finite system size can influence the long-term system dynamics by inducing spontaneous transitions.
+These results provide important insights and can have diverse applications, such as designing optimal surgical procedures for drug-resistant epilepsy in the human brain and identifying the sources of rumours in social networks.
+keywords:
+Higher-order interactions, adaptive connectivity, explosive phase transitions, synchronization, epilepsy.
+journal: Chaos, Solitons & Fractals
+[…]
+[label1]organization=National Brain Research Centre,addressline=Manesar, city=Gurgaon, postcode=122052, state=Haryana, country=India
+1 Introduction
+Recent literature across different scientific disciplines shows that real-life complex physical systems involve connections that extend beyond basic pairwise edges and include couplings that capture interactions between groups of three or more nodes simultaneously [hypergraphtheoryBretto2013, majhi2022dynamics, shi2022simplicial, boccaletti2023structure, PhysRevLett.131.207401]; see Fig.
 1(a) for a sample schematic.
-Interactions in social dynamics are an archetypal example and include collaborations in scientific research [lungeanu2021team, moore2012analyzing], group interactions in opinion spreading [wang2020social, ghahremani2023novel], and dynamics of consensus [PhysRevE.101.032310, PhysRevE.104.064305], among others.
-In biological systems, which are often replete with higher-order interactions, neuroscience stands out as a promising domain [giusti2016two, hindriks2025unraveling].
+Such interactions, mathematically represented as hyper-edges or simplicial complexes [salnikov2018simplicial], cannot be further decomposed into their constituent pairwise interactions.
+Systems arising from social dynamics are an archetypal example, and include collaboration networks in scientific research [lungeanu2021team, moore2012analyzing], group interactions in opinion spreading [wang2020social, ghahremani2023novel], and dynamics of consensus [PhysRevE.101.032310, PhysRevE.104.064305], among others.
+Within biological systems, which are often replete with higher-order interactions, neuroscience stands out as a promising domain [giusti2016two, hindriks2025unraveling].
 Higher-order connectivity is evident in the human brain [karmelic2022emergent] (also see Fig.
@@
 These also appear in models describing chemical reactions [wen2023chemical, brauch2013higher], power grids [ghasemi2022higher, PhysRevE.109.024212], and the evolution of ecological systems [swain2022higher, gallardo2024higher, bairey2016high, PhysRevE.111.014309], highlighting their ubiquity across domains.
-Figure 1: (a): Schematic representation of hyperedges in the brain: In (i), pink circles indicate nodes, or Broadman areas of functional significance, whereas the lines (edges) and coloured regions (hyperedges) denote pairwise and higher-order interactions, respectively.
-These can appear at multiple spatial scales, such as macroscopic functional connectivity patterns from fMRI-BOLD time series and microscopic neuronal scale, shown in (ii) and (iii), respectively; (b): Number of interactions in a brain network, as a function of its order, relative to pairwise interactions [karmelic2022emergent, cabral2022metastable]; (c): Temporal variation of intracranial (iEEG) signals from different channels [bougou2025mesoscale, dataset], with the histogram (top-inset) and time series (bottom-inset) of coherence, highlighting its increase during epilepsy, the onset of which is marked with a vertical red line.
-The dynamics of systems with higher-order interactions can be described by equations of the form
+Mathematically, the dynamics of such systems can be described by
 $\frac{{\rm d}x_{i}}{{\rm d}t}$ $=f_{i}(x_{i})+\sum_{\mathclap{d=1;j_{1},..,j_{d}=1}}^{D;N}A^{(d)}_{i,j_{1},..,j_{d}}g^{(d)}(x_{i},x_{j_{1}},..,x_{j_{d}}),$ (1)
 where $N$ and $D$ denote the number of constituent nodes and the maximum order of interaction, respectively.
-Oscillations are a characteristic feature often observed in such systems [rabinovich2012oscillations], with their synchronization representing a vital emergent phenomenon having important implications across disciplines [arenas2008synchronization]; for example, epilepsy is often characterized by synchronous activity in the brain [osterhage2007measuring, lehnertz2009synchronization]; see Fig.
-1(c) for details.
-To this end, the Kuramoto model of coupled phase oscillators [RevModPhys.77.137, gupta2014kuramoto] serves as a paradigmatic framework for quantitatively exploring synchronization in a systemic manner.
-To capture realistic systems, it has also been extensively modified with features such as complex connectivity matrices [jain2023composed, chen2023critical, peron2019onset, tong2018exponential], symmetry-breaking mechanisms [manoranjani2023generalization, biswas2024symmetry], time-delayed interactions [yeung1999time, wu2018dynamics, ross2021dynamics, budzinski2023analytical], higher-order couplings [xu2021spectrum, costa2024bifurcations, costa2025exact, moyal2024rotating, rajwani2023tiered, carballosa2023cluster, sabhahit2024prolonged, skardal2020higher], as well as different combinations of each of these [skardal2022tiered].
-These find applications in a myriad of different domains [guo2021overviews, wang2021dynamic, blanter2016kuramoto, fioriti2008stability], and are of particular importance in neuroscience, where they provide a tractable mathematical framework to model complex neuro-oscillatory phenomena [odor2019critical, heggli2019kuramoto, PhysRevLett.120.244101, odor2021effect, pathak2022biophysical, ranjan2024propagation, PhysRevE.111.044310].
-This work investigates an adaptive Kuramoto model, which incorporates arbitrary-order interactions of both symmetric and asymmetric types, to explore the effects of higher-order interactions and their interplay with synchronization.
+Here, the tensor $A^{(d)}$ and the multivariate function $g^{(d)}(\cdot)$ capture the structure and mechanism of interactions, respectively, of order $d\leq D$.
+Oscillations are a characteristic feature often observed in such systems, with their synchronisation representing a vital emergent phenomenon having important implications across disciplines [arenas2008synchronization].
+It is particularly important in models of regular and pathological neural dynamics [ansarinasab2025synchronization, ansarinasab2025transition], among which a key example is epilepsy, often characterised by excess synchronous activity between different regions of the brain [osterhage2007measuring, lehnertz2009synchronization]; see Fig.
+1(c) for sample schematics.
+To this end, the Kuramoto model [RevModPhys.77.137, gupta2014kuramoto] serves as a paradigmatic framework for quantitatively exploring synchronisation in a systematic manner.
+In addition, to capture physically realistic complex systems, it has also been extensively modified to include features such as complex connectivity patterns [jain2023composed, chen2023critical, peron2019onset, tong2018exponential], symmetry-breaking mechanisms [PhysRevE.102.012206, manoranjani2023generalization, biswas2024symmetry], time-delayed interactions [yeung1999time, wu2018dynamics, ross2021dynamics, skardal2022tiered, budzinski2023analytical], phase frustration [buzanello2022matrix, PhysRevE.108.034208, moyal2024rotating, mondal2025enhancing], and higher-order couplings [xu2021spectrum, costa2024bifurcations, costa2025exact, rajwani2023tiered, carballosa2023cluster, sabhahit2024prolonged, skardal2020higher], as well as different combinations of each of these.
+The Kuramoto model finds applications in a myriad of different domains [guo2021overviews, wang2021dynamic, blanter2016kuramoto, fioriti2008stability], and is of particular importance in neuroscience, where it provides a tractable mathematical framework to capture complex neuro-oscillatory phenomena [odor2019critical, heggli2019kuramoto, PhysRevLett.120.244101, odor2021effect, pathak2022biophysical, ranjan2024propagation, PhysRevE.111.044310].
+Additionally, it has also been extended to the so-called “swarmalator” model [o2017oscillators, PhysRevE.111.014209, ansarinasab2024spatial], which introduces additional state variables to capture the dynamics of agents moving through space and interacting through a combination of both phase and spatial interactions, thereby demonstrating a rich repertoire of dynamical states.
+Figure 1: (a): Schematic representation of hyperedges in the brain: In (i), pink circles indicate nodes, or Broadman areas of functional significance, whereas the lines (edges) and coloured regions (hyperedges) denote pairwise and higher-order interactions, respectively.
+These can appear at multiple spatial scales, such as macroscopic functional connectivity patterns from fMRI-BOLD time series and microscopic neuronal scale, shown in (ii) and (iii), respectively; (b): Number of interactions in a brain network, as a function of its order, relative to pairwise interactions; see Refs.
+[karmelic2022emergent, cabral2022metastable] for more details; (c): Variation of intracranial (iEEG) signals from different channels [bougou2025mesoscale, dataset], with the histogram (top-inset) and time series (bottom-inset) of phase synchrony, highlighting its increase during epilepsy, with its onset marked with a vertical line.
+To gain an understanding of synchronisation phenomena, this article investigates an adaptive Kuramoto model that incorporates arbitrary-order interactions of both symmetry-preserving and symmetry-breaking types to explore the effects of higher-order interactions and their interplay with other coupling modalities.
 The dynamics of the system are assumed to be governed by equations of the form
 $\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r_{1})\epsilon_{1}}{N}\sum_{j_{1}=1}^{N}\sin\Big(\phi_{j_{1}}-\phi_{i}-\delta_{1}\Big)+\\ \sum_{d=2}^{d=D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{N^{d}}\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(d\phi_{j_{d}}-\sum_{k=1}^{d-1}\phi_{j_{k}}-\phi_{i}-\delta_{d}\right)+\right.\\ \left.\frac{f_{d}^{\prime}(r_{1})\epsilon^{\prime}_{d}}{N^{d}}\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(d\phi_{j_{d}}+\sum_{k=1}^{d-1}\phi_{j_{k}}-\phi_{i}-\delta^{\prime}_{d}\right)\right],$ (2)
-where $z_{k}=r_{k}e^{{\rm i}\psi_{k}}=N^{-1}\sum_{j=1}^{N}e^{{\rm i}k\phi_{j}}\ \forall\ k\geq 1$ are the Kuramoto-Daido order parameters [clusella2020irregular] and ${\rm i}=\sqrt{-1}$.
-In this model, the interaction terms are strongly influenced by the corresponding adaptation functions, the inclusion of which represents a fundamental dynamical feature of natural systems [sawicki2023perspectives].
-Adaptation allows the characteristic properties of the system to change in response to the collective state of the system and is seen across bio-physical, socio-economic, and engineering systems, and is particularly prevalent in the study of neural activity; for example, spike-frequency adaptation models the reduction in neural firing arising from the fatigue of continuous spiking activity.
-[PhysRevE.107.024311].
-In the context of synchronization, this allows the modulation of the global coupling strength in response to the extent of synchronization in the system [jin2023synchronization], thereby allowing the system to reach a desired state of synchronicity.
-A universal and well-investigated form of the adaptation function is a power-law, given by $f(r_{1})\propto r_{1}^{\alpha}$, where $\alpha$ controls the nature of the feedback and the resulting dynamics [PhysRevE.75.017201, PhysRevE.102.012219, biswas2024effect].
-Analysis — For an infinitely large number of oscillators, an assumption relaxed later, the state of the system can be described by a probability density function $F(\phi,\omega;t)$, which satisfies a continuity equation of the form $\partial_{t}F+\partial_{\phi}(F\dot{\phi})=0$.
-Since $F(\phi+2\pi,\omega;t)=F(\phi,\omega;t)$, it can be expanded in a Fourier series in $\phi$, where $F_{n}(\omega,t)$ is the $n^{\rm th}$ Fourier coefficient and $F_{-n}(\omega,t)=F^{*}_{n}(\omega,t)$.
-Using the Ott-Antonsen ansatz [ott2008low, ott2009long], it can be shown that
-$\frac{\partial v}{\partial t}+{\rm i}\omega v+\frac{f_{1}(r_{1})\epsilon_{1}}{2}\left(z_{1}e^{-{\rm i}\delta_{1}}v^{2}-z_{1}^{*}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}\delta_{d}}v^{2}-z_{d}^{*}z_{1}^{d-1}e^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}\delta^{\prime}_{d}}v^{2}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}\delta^{\prime}_{d}}\right)\right]=0,$ (3)
-where $F_{n}(\omega,t)=v^{n}(\omega,t)$.
-Subsequently, by assuming that $\omega_{i}$ are sampled from a Cauchy distribution with mean $\omega_{0}$ and scale factor $\Delta$, the order parameters are obtained as $z_{d}=\left[v^{*}(\omega_{0}-i\Delta,t)\right]^{d}$.
-Substituting into Eq.
-(3) and re-labelling $z_{1}$ as $z$ for simplicity, it takes the form
-$\frac{\partial z}{\partial t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{f_{1}(r)\epsilon_{1}}{2}\left(ze^{-{\rm i}\delta_{1}}-z|z|^{2}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r)\epsilon_{d}}{2}\left(|z|^{2(d-1)}ze^{-{\rm i}\delta_{d}}-|z|^{2d}ze^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r)\epsilon^{\prime}_{d}}{2}\left(z^{2d-1}e^{-{\rm i}\delta^{\prime}_{d}}-|z|^{4}(z^{*})^{2d-3}e^{{\rm i}\delta^{\prime}_{d}}\right)\right].$ (4)
-Evidently, the incoherent state is stable only for $\epsilon_{1}\leq\epsilon_{1}^{c}=2\Delta/f_{1}(0)\cos(\delta_{1})$, which denotes the transition point from an asynchronous state to a synchronous state.
-In contrast, the dynamics of the synchronous states are governed by equations of the form
+where $\phi_{i}(t)$ and $\omega_{i}$ represent the instantaneous phase and natural frequency, respectively, of each oscillator $i$. Furthermore,
+$z_{k}=r_{k}e^{{\rm i}\psi_{k}}=\frac{1}{N}\sum_{j=1}^{N}e^{{\rm i}k\phi_{j}},\quad k\geq 1,$ (3)
+are the complex Kuramoto-Daido order parameters [clusella2020irregular] that quantify the collective dynamics of the systems.
+In particular, for a nearly incoherent state, $r_{1}$ tends to zero, whereas for a highly synchronous state, $r_{1}$ approaches unity.
+In this model, the couplings between the individual phases are captured through the second, third, and fourth terms in the RHS of Eq.
+(2), which represent the pairwise and the two separate types of higher-order interactions, respectively.
+Evidently, the system is globally coupled, i.e., every node interacts with every other node in an identical manner, across both pairwise and higher-order links.
+The interaction terms are strongly influenced by the functions $f_{d}(\cdot)$ and $f_{d}^{\prime}(\cdot)$, the inclusion of which captures adaptivity, a fundamental feature of many natural systems [sawicki2023perspectives], that allows modulating the characteristic properties in response to the current collective state of the overall system.
+It is seen across bio-physical, socio-economic, and engineering systems, and is particularly prevalent in the study of neural activity; for example, spike-frequency adaptation models the reduction in neural firing arising from the fatigue of continuous spiking activity [PhysRevE.107.024311].
+In the context of synchronisation, this allows the modulation of the global coupling strengths (i.e., $\epsilon_{d}$ and $\epsilon_{d}^{\prime}$) in response to the extent of synchronisation in the system [jin2023synchronization], quantified through $r_{1}$, thereby allowing the system to reach a desired state of synchronicity.
+For example, a universal and well-investigated form of the adaptation function is a power-law, where the exponent controls the nature of the feedback and the resulting dynamics [PhysRevE.75.017201, PhysRevE.102.012219, biswas2024effect], a variation of which is investigated in later sections.
+Furthermore, the coupling terms also comprise of phase lags in each order of interaction (i.e., $\delta_{d}$ and $\delta_{d}^{\prime}$), the primary effect of which is to modify the common frequency to some desired value [lohe2015synchronization], whereas both symmetry-preserving and symmetry-breaking higher-order interactions terms are introduced to capture non-trivial generalizations of the basic Kuramoto model, such as the Winfree model [manoranjani2023generalization, PhysRevE.96.042208].
+The rest of the article is arranged as follows: Sec.
+2 outlines the dimensionality reduction of the proposed system, which provides analytical insights into the overall dynamics.
+In Sec.
+3, these results are validated against numerical simulations in Sec.
+3.1, highlighting different dynamical regimes, with subsequent subsections showcasing the relative effect of interactions (Sec.
+3.2) and the effect of fluctuations in finite-sized networks (Sec.
+3.3).
+The paper ends with a summary of the key results, limitations, and future research directions in Sec.
+4.
+2 Analytical Insights
+To obtain analytical insights into the dynamics of the proposed system, Eq. (2) is first rewritten, using the definition of the order parameter $z_{k}$ in Eq. (3), in the form
+$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r_{1})\epsilon_{1}}{2{\rm i}}\left(z_{1}e^{-{\rm i}(\delta_{1}+\phi_{i})}-z_{1}^{*}e^{{\rm i}(\delta_{1}+\phi_{i})}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2{\rm i}}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}(\delta_{d}+\phi_{i})}-z^{*}_{d}z_{1}^{d-1}e^{{\rm i}(\delta_{d}+\phi_{i})}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2{\rm i}}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}(\delta^{\prime}_{d}+\phi_{i})}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}(\delta^{\prime}_{d}+\phi_{i})}\right)\right],$ (4)
+where * represents complex conjugation.
+For an infinitely large number of oscillators, an assumption which is relaxed in the later sections, the state of the system in Eq.
+(2) can be described by a joint probability density function $F(\phi,\omega;t)$, which satisfies a continuity equation of the form
+$\frac{\partial F}{\partial t}+\frac{\partial(F\dot{\phi})}{\partial\phi}=0$ (5)
+and the condition
+$\int_{0}^{2\pi}F(\phi,\omega;t)\ {\rm d}\phi=g(\omega),$ (6)
+where $\dot{\phi}$ is given by Eq. (4). Since $F(\phi+2\pi,\omega;t)=F(\phi,\omega;t)$, it can also be expanded as a series of the form
+$F(\phi,\omega;t)=\frac{g(\omega)}{2\pi}\left[1+\sum_{n=1}^{\infty}\left(F_{n}e^{{\rm i}n\phi}+F^{*}_{n}e^{-{\rm i}n\phi}\right)\right],$ (7)
+where $F_{n}\equiv F_{n}(\omega,t)$ is the $n^{\rm th}$ Fourier coefficient.
+By using the Ott-Antonsen ansatz [ott2008low, ott2009long], the Fourier modes are assumed to be of the form $F_{n}(\omega,t)=v^{n}(\omega,t)$, with the restriction $|v(\omega,t)|<1$ so that the sum in Eq.
+(7) converges.
+This implies that Eq.
+(5) can be re-written in terms of $v(\omega,t)$ as
+$\frac{\partial v}{\partial t}+i\omega v+\frac{f_{1}(r_{1})\epsilon_{1}}{2}\left(z_{1}e^{-{\rm i}\delta_{1}}v^{2}-z_{1}^{*}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r_{1})\epsilon_{d}}{2}\left(z_{d}(z_{1}^{*})^{d-1}e^{-{\rm i}\delta_{d}}v^{2}-z_{d}^{*}z_{1}^{d-1}e^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r_{1})\epsilon_{d}^{\prime}}{2}\left(z_{d}z_{1}^{d-1}e^{-{\rm i}\delta^{\prime}_{d}}v^{2}-z_{d}^{*}(z_{1}^{*})^{d-1}e^{{\rm i}\delta^{\prime}_{d}}\right)\right]=0.$ (8)
+Assuming that the natural frequencies $\omega_{i}$ are sampled from a Cauchy distribution with mean $\omega_{0}$ and scale factor $\Delta$, the summation in Eq.
+(3) can be replaced with an integral, which can then be evaluated to obtain
+$z_{d}=\int_{0}^{2\pi}\int_{-\infty}^{\infty}e^{{\rm i}d\phi}F(\phi,\omega;t)\ {\rm d}\omega\ {\rm d}\phi=\Big[v^{*}(\omega_{0}-i\Delta,t)\Big]^{d}.$ (9)
+Upon by substituting Eq. (9) into Eq. (8) and re-labelling $z_{1}$ as $z$ for the sake of simpler notation, Eq. (8) takes the form
+$\frac{{\rm d}z}{{\rm d}t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{f_{1}(r)\epsilon_{1}}{2}\left(ze^{-{\rm i}\delta_{1}}-z|z|^{2}e^{{\rm i}\delta_{1}}\right)+\\ \sum_{d=2}^{D}\left[\frac{f_{d}(r)\epsilon_{d}}{2}\left(|z|^{2(d-1)}ze^{-{\rm i}\delta_{d}}-|z|^{2d}ze^{{\rm i}\delta_{d}}\right)\right.+\\ \left.\frac{f^{\prime}_{d}(r)\epsilon^{\prime}_{d}}{2}\left(z^{2d-1}e^{-{\rm i}\delta^{\prime}_{d}}-|z|^{4}(z^{*})^{2d-3}e^{{\rm i}\delta^{\prime}_{d}}\right)\right],$ (10)
+yielding a two-dimensional coupled system, the analysis of which is simpler compared to Eq.
+(2).
+First, the stability of the incoherent state of the system can be obtained by linearizing Eq.
+(10) about $z(t)=0$, leading to the equation
+$\frac{{\rm d}z}{{\rm d}t}={\rm i}(\omega_{0}+{\rm i}\Delta)z+\frac{z}{2}f_{1}(0)\epsilon_{1}e^{-{\rm i}\delta_{1}}.$ (11)
+This implies that $z=0$ is stable only for
+$\epsilon_{1}\leq\epsilon_{1}^{c}=\frac{2\Delta}{f_{1}(0)\cos(\delta_{1})},$ (12)
+which corresponds to the critical point beyond which the system undergoes a phase transition from an asynchronous state to a synchronous state. Alternatively, Eq. (10) can also be rewritten as
+$\begin{split}\frac{{\rm d}r}{{\rm d}t}+{\rm i}r\frac{{\rm d}\psi}{{\rm d}t}&={\rm i}(\omega_{0}+{\rm i}\Delta)r+\frac{rf_{1}(r)\epsilon_{1}}{2}H(r,\delta_{1})+\\ &\sum_{d=2}^{D}r^{2d-1}\left[\frac{f_{d}(r)\epsilon_{d}}{2}H(r,\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}H(r,\chi^{\prime}_{d})\right],\end{split}$ (13)
+where $H(x,y)=(1-x^{2})\cos(y)-{\rm i}(1+x^{2})\sin(y)$ and $\chi^{\prime}_{d}=\delta^{\prime}_{d}-2(d-1)\psi$.
+Therefore, the dynamics of the synchronous states (i.e., where $r>0$) are governed by coupled differential equations of the form
 $\frac{{\rm d}r}{{\rm d}t}=R(r,\psi)=-\Delta r+r(1-r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\cos(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\cos(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\cos(\chi^{\prime}_{d})\right)\right],$ (14)
 and
-$\frac{{\rm d}\psi}{{\rm d}t}=S(r,\psi)=\omega_{0}-(1+r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\sin(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\sin(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\sin(\chi^{\prime}_{d})\right)\right],$ (6)
-where $\chi^{\prime}_{d}=\delta^{\prime}_{d}-2(d-1)\psi$.
-These two equations dictate the dynamics of $r(t)$ and $\psi(t)$, the steady state solution $(r_{s},\psi_{s})$, and the subsequent stability analysis provides insights into the overall system dynamics.
+$\frac{{\rm d}\psi}{{\rm d}t}=\Psi(r,\psi)=\omega_{0}-(1+r^{2})\left[\frac{f_{1}(r)\epsilon_{1}}{2}\sin(\delta_{1})\right.+\\ \left.\sum_{d=2}^{D}r^{2(d-1)}\left(\frac{f_{d}(r)\epsilon_{d}}{2}\sin(\delta_{d})+\frac{f^{\prime}_{d}(r)\epsilon_{d}^{\prime}}{2}\sin(\chi^{\prime}_{d})\right)\right],$ (15)
+which can be obtained by separating the real and imaginary parts of Eq.
+(13).
+The steady state solutions of the above two equations (say, $(r_{s},\psi_{s})$) provide insights into the system dynamics.
+These can be obtained as solutions of the algebraic equations $R(r_{s},\psi_{s})=0$ and $\Psi(r_{s},\psi_{s})=0$, which can be expressed as
+$\left(\frac{\Delta}{1-r^{2}_{s}}\right)-\sum_{n=1}^{D}\sigma_{n}(r_{s})\cos(\delta_{n})=\sum_{m=2}^{D}\sigma_{m}^{\prime}(r_{s})\cos(\chi_{m,s}^{\prime}),$ (16)
+and
+$\left(\frac{\omega_{0}}{1+r^{2}_{s}}\right)-\sum_{n=1}^{D}\sigma_{n}(r_{s})\sin(\delta_{n})=\sum_{m=2}^{D}\sigma_{m}^{\prime}(r_{s})\sin(\chi_{m,s}^{\prime}),$ (17)
+where $\sigma_{n}(r)=f_{n}(r)\epsilon_{n}r^{2(n-1)}/2$, $\sigma^{\prime}_{n}(r)=f^{\prime}_{n}(r)\epsilon^{\prime}_{n}r^{2(n-1)}/2$, and $\chi^{\prime}_{m,s}$ denotes $\chi^{\prime}_{m}$ evaluated at $\psi={\psi_{s}}$.
+Obtaining their stability entails calculating the Jacobian $J$ of Eqs.
+(14) and (15), the analytical expressions of its elements $J_{ij}$ are listed in Eqs.
+(22)-(25) of A.
+Therefore, taken together, the analytical results presented in this section provide a theory of synchronization in systems with arbitrary-order interactions, while also accounting for the effects of different adaptive mechanisms, phase lags, and generalized forms of interactions that break rotational symmetry in the Kuramoto model.
+The subsequent sections are devoted to verifying these results through numerical simulations, as well as investigating other properties of the proposed system through numerical simulations.
+3 Numerical Results
+Although the analytical results derived in the previous section constitute a general, arbitrary-order theoretical framework, further analytical progress is feasible only under specific simplifying assumptions on the parameter values, whereas incorporating arbitrarily large orders of interactions would increase simulation times beyond reasonable limits.
+Therefore, to validate the obtained analytical expressions, the simplest yet representative example of Eq.
+(2) is chosen.
+It corresponds to a system consisting of pairwise and triadic interactions, similar to that reported in Ref.
+[biswas2024symmetry], and is described by equations of the form
+$\begin{split}\frac{{\rm d}\phi_{i}}{{\rm d}t}=&\omega_{i}+\frac{f_{1}(r)\epsilon_{1}}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i})+\\ &\frac{f_{2}(r)\epsilon_{2}}{N^{2}}\sum_{j,k=1}^{N}\Bigg[\sin(2\phi_{k}-\phi_{j}-\phi_{i})+\xi\sin(2\phi_{k}+\phi_{j}-\phi_{i})\Bigg],\end{split}$ (18)
+where $\xi=\epsilon_{2}^{\prime}/\epsilon_{2}$.
+For the sake of simplicity, $f_{d}(r)$ are assumed to be of the form $f_{d}(r)=(1+\gamma r)^{\pm 2}$ for $d=1$ and $d=2$, respectively, whereas $\gamma$ is a constant that generates a family of qualitatively different functional forms; also see Fig.
+2.
+(a) $d=1$
+(b) $d=2$
+Figure 2: Plot of $f_{d}(r)$ as a function of $r\in[0,1]$ for (a) $d=1$ and (b) $d=2$, where the colours denote different values of $\gamma$.
+This choice represents the simplest possible functional form of a polynomial adaptation function, which has previously been shown to induce both continuous and explosive phase transitions [biswas2024effect], and has been set up such that it introduces qualitatively different adaptation behaviour (i.e., increasing and decreasing) across the two orders of interactions.
 (a) $\epsilon_{1}=0.6$
@@
 For all figures, $\omega_{0}=1.5$, $\gamma=0.2$, and $\epsilon_{2}=5.0$.
-Results — As a relatively simple example of Eq. (2), a system described by equations of the form
-$\frac{{\rm d}\phi_{i}}{{\rm d}t}=\omega_{i}+\frac{f_{1}(r)\epsilon_{1}}{N}\sum_{j=1}^{N}\sin(\phi_{j}-\phi_{i})+\\ \frac{f_{2}(r)\epsilon_{2}}{N^{2}}\sum_{j,k=1}^{N}\Bigg[\sin(2\phi_{k}-\phi_{j}-\phi_{i})+\xi\sin(2\phi_{k}+\phi_{j}-\phi_{i})\Bigg],$ (7)
-is investigated, where $\xi=\epsilon_{2}^{\prime}/\epsilon_{2}$.
-For the sake of simplicity, $f_{d}(r)$ are assumed to be of the form $f_{d}(r)=(1+\gamma r)^{\pm 2}$ for $d=1$ and $d=2$, respectively, whereas $\gamma$ is a constant.
-In this case, Eqs.
-(5) and (6) take the form
-$\dot{r}$ $=-\Delta r+r(1-r^{2})\Big[\sigma_{1}(r)+\sigma_{2}(r)\Big(1+\xi\cos(2\psi)\Big)\Big],$ (8)
-$\dot{\psi}$ $=\omega_{0}+\xi(1+r^{2})\sigma_{2}(r)\sin(2\psi),$ (9)
-respectively, where $\sigma_{n}(r)=f_{n}(r)\epsilon_{n}r^{2(n-1)}/2$ for $n=1,2$, and $\epsilon_{1}^{c}=2\Delta$. The value of $r_{s}$ can be obtained as a solution of the algebraic equation
-$\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right]^{2}+\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right]^{2}=\xi^{2}\sigma_{2}^{2}(r_{s}).$ (10)
-However, in most cases, it is not possible to find closed-form analytical solutions to the above equations, thus requiring the use of numerical techniques.
+In this case, Eqs. (14) and (15) take the form
+$\frac{{\rm d}r}{{\rm d}t}$ $=-\Delta r+r(1-r^{2})\Big[\sigma_{1}(r)+\sigma_{2}(r)\Big(1+\xi\cos(2\psi)\Big)\Big],$ (19)
+and
+$\frac{{\rm d}\psi}{{\rm d}t}$ $=\omega_{0}+\xi(1+r^{2})\sigma_{2}(r)\sin(2\psi),$ (20)
+respectively, where the forward phase transition point is obtained as $\epsilon_{1}^{c}=2\Delta$ from Eq.
+(12).
+Therefore, the value of $r$ in equilibrium (i.e., $r_{s}$) can be obtained as a solution of the algebraic equation
+$\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right]^{2}+\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right]^{2}=\xi^{2}\sigma_{2}^{2}(r_{s}),$ (21)
+the stability of which is obtained from the Jacobian of Eqs.
+(19) and (20); see Eqs.
+(26)-(29) in A.
+However, finding closed-form analytical solutions is not always feasible, thus requiring the use of numerical techniques.
+Furthermore, even while limited to triadic interactions, the computational load for simulating Eq.
+(18) is extensive due to the presence of nested summations in the coupling term, especially for large system sizes.
+This necessitates the development of alternative schemes that reduce the computational burden, an example of which is outlined in B.
+3.1 Characteristic system dynamics
 For the rest of the work, $(\Delta,\xi)=(1.0,0.4)$ in Eq.
 (18) for the sake of simplicity.
-Furthermore, as an example of adaptive systems, the specific case of $\gamma=0.2$ is demonstrated in Fig.
-2, for different values of $\epsilon_{1}$, whereas the value of $\omega_{0}$ and $\epsilon_{2}$ is set at $1.5$ and $5.0$, respectively  [^1] The non-adaptive case (i.e., $\gamma=0$) has already been studied in Ref.
-[biswas2024symmetry]..
-For $\epsilon_{1}=0.6$, it is evident that Eq.
+Furthermore, as an example of adaptive systems [^1] The non-adaptive case (i.e., $\gamma=0$) has already been studied in Ref.
+[biswas2024symmetry], facilitating direct comparison., the specific case of $\gamma=0.2$ is demonstrated in Fig.
+3, for different values of $\epsilon_{1}$, whereas the value of $\omega_{0}$ and $\epsilon_{2}$ is set at $1.5$ and $5.0$, respectively.
+In this scenario, for $\epsilon_{1}=0.6$, it is evident that Eq.
 (21) admits no solutions.
@@
 Thus, the system always converges towards $r(t)\approx 0$, irrespective of initial conditions; see Fig.
-2(a).
+3(a).
 However, as the value of $\epsilon_{1}$ is increased to $1.3$, there is a qualitative change in the dynamics.
 In this case, the asynchronous state remains stable, since $\epsilon_{1}<\epsilon_{1}^{c}$, and is evident from the temporal variation of $r(t)$, which converges towards $r(t)\approx 0$ when initialised in an asynchronous state.
-However, the system also demonstrates an unsteady synchronised state, where the order parameter shows oscillations about a mean value; see Fig.
-2(b).
+However, the system also demonstrates an unsteady synchronised state, where the order parameter shows temporal oscillations as highlighted in Fig.
+3(b), a behaviour previously observed in Ref.
+[biswas2024symmetry] in a similar model, and also in Ref.
+[biswas2022mirroring] where it emerges due to the formation of partial clusters of oscillators which instantaneously overlap and separate.
 (a) $\gamma=0.05$
@@
 (j) $\epsilon_{2}=6.0$
-Figure 3: (a)-(j): Variation of steady state order parameter $r$ as a function of $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ for different values of $\gamma$ (top-row; $\epsilon_{2}=5.0$), and $\epsilon_{2}$ (bottom-row; $\gamma=0.2$).
-The vertical line corresponds to the forward transition point (i.e., $E^{c}=\epsilon_{1}^{c}/\epsilon_{2}$), whereas the blue and red curves denote the forward and backward variation of $E$, respectively.
-As demonstrated in Fig.
-2(c), increasing $\epsilon_{1}$ to $1.8$ entails further change in the dynamics.
+Figure 4: (a)-(j): Variation of the order parameter $r$ as a function of $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ for different values of $\gamma$ (top-row; $\epsilon_{2}=5.0$), and $\epsilon_{2}$ (bottom-row; $\gamma=0.2$).
+The vertical line corresponds to the forward transition point (i.e., $E^{c}=\epsilon_{1}^{c}/\epsilon_{2}$), whereas the blue and red curves denote the forward and backward variation of $E$, respectively, whereas $\omega_{0}=1.5$ across all figures.
+The observations in Fig.
+3(c) demonstrate that increasing $\epsilon_{1}$ to $1.8$ entails further change in the dynamics.
 In this case, Eq.
@@
 In particular, the asynchronous initial conditions continue to be asynchronous, whereas the synchronous initial condition converges to $r(t)\approx 0.82$.
-Upon further increasing $\epsilon_{1}$ to $2.6$, the asynchronous state loses stability since $\epsilon_{1}>\epsilon_{1}^{c}$, as demonstrated in Fig.
-2(d).
+Finally, upon further increasing $\epsilon_{1}$ to $2.6$, the asynchronous state loses stability since $\epsilon_{1}>\epsilon_{1}^{c}$, as demonstrated in Fig.
+3(d).
 However, Eq.
@@
 Thus, the system is now monostable, which is demonstrated by the temporal variation of $r(t)$, which converges around $r(t)\approx 0.82$ irrespective of the initial state of the system.
-Relative effect of interactions: To quantify the relative effect of pairwise and higher-order interactions, the dynamics of the system is investigated via the variation in the steady state value of the order parameter as a function of the ratio $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ in Fig.
-3, representing the relative strength of the pairwise and triadic interactions.
+3.2 Relative effect of interactions
+To quantify the relative effect of pairwise and higher-order interactions, the dynamics of the system are investigated via the variation in the steady state value of the order parameter as a function of the ratio $E=\epsilon_{1}/\epsilon_{2}\in[0,2]$ in Fig.
+4, representing the relative strength of the pairwise and triadic interactions.
 The top and bottom rows correspond to $\epsilon_{2}=5.0$ and $\gamma=0.2$, respectively.
-In the former, the value of $\gamma$ varies in the range $[0.05,0.25]$ in steps of $0.05$ across the sub-figures, whereas in the latter, the value of $\epsilon_{2}$ varies in the range $[2.0,6.0]$ in steps of $1.0$.
+In the former, the value of $\gamma$ varies between $[0.05,0.25]$ in steps of $0.05$ across the sub-figures, whereas in the latter, the value of $\epsilon_{2}$ varies between $[2.0,6.0]$ in steps of $1.0$.
 In Figs.
-3(a)-3(e), upon increasing $\gamma$, the transition to synchrony upon forward variation of $E$ is observed to be explosive, which is characterized by a sudden increase in $r$ at the critical point $E\approx 0.4$.
+4(a)-4(e), upon increasing $\gamma$, the transition to synchrony upon forward variation of $E$ is observed to be explosive, which is characterised by a sudden increase in $r$ at the critical point $E\approx 0.4$.
 In contrast, the backward transition point, captured by the sharp drop in $r$ upon reducing $E$, reduces with increasing values of $\gamma$.
 However, the post-transition state upon increasing $E$, as well as the route to desynchronization upon reducing $E$, varies for different values of $\gamma$.
-In particular, for $\gamma=0.05$, the system demonstrates an unsteady synchronous state characterized by an oscillatory temporal variation of $r(t)$ immediately after the explosive transition.
+In particular, for $\gamma=0.05$, the system demonstrates an unsteady synchronous state characterised by an oscillatory temporal variation of $r(t)$ immediately after the explosive transition.
 Furthermore, tuning $E$ beyond $1.0$ results in a steady synchronous state; see Fig.
-3(a).
+4(a).
 Subsequently, upon increasing $\gamma$ to $0.10$ and $0.15$, the region of $E$ for which the system demonstrates an unsteady synchronous state reduces, resulting in a larger parameter range for which the steady synchronous state is observed; see Figs.
-3(b) and 3(c).
+4(b) and 4(c).
 Furthermore, upon increasing $\gamma$ to $0.20$ and $0.25$, the system exhibits only a steady synchronous state after the transition.
 However, a key difference is that, in the former, the system exhibits an unsteady synchronous state in a portion of the bi-stable regime before desynchronization, whereas the latter maintains the steady synchronous state en route to desynchronization; see Figs.
-3(d) and 3(e).
+4(d) and 4(e).
 The curves in Figs.
-3(f)-3(j) demonstrate the qualitative dynamical changes upon changing $\epsilon_{2}$.
+4(f)-4(j) demonstrate the qualitative dynamical changes upon changing $\epsilon_{2}$.
 Unlike the previous case, changing $\epsilon_{2}$ modifies the characteristic nature of the transition to synchrony upon varying $E$.
-In particular, while for $\epsilon_{2}=2.0$, the transition is observed to be continuos with a smooth increase in the steady-state value of $r$ beyond the critical point (see Fig.
-3(f)), higher values of $\epsilon_{2}$ result in explosive synchronization transitions, in addition to the formation of a bi-stable zone; see Figs.
-3(g)-3(j).
+In particular, while for $\epsilon_{2}=2.0$, the transition is observed to be continuous with a smooth increase in the steady-state value of $r$ beyond the critical point (see Fig.
+4(f)), higher values of $\epsilon_{2}$ result in explosive synchronization transitions, in addition to the formation of a bi-stable zone; see Figs.
+4(g)-4(j).
 Similar to the previous case, the post-transition synchronous states can be both unsteady or steady (for example, see Figs.
-3(h) and 3(i)), whereas, for $\epsilon_{2}=6.0$, the system does not undergo any de-synchronisation transition upon reducing $E$; see Fig.
-3(j).
-Effect of system size: Real bio-physical systems, such as the brain, are typically modelled as a connected system with a finite number of constituent dynamical units (here, oscillators), the size of which could be of different orders depending on the level of abstraction.
-Thus, while the previous sections demonstrate steady-state results for an infinitely large number of oscillators, the synchronization phenomena in finite-sized systems are typically affected by fluctuations which can be better characterised by constructing a probability density of the order parameter, denoted by $p(r)$.
-This is computed by simulating the system for a large number of time-steps and repeated over multiple initial conditions, as well as for two qualitatively different sets characterised by $r(0)\approx 0.0$ (i.e., nearly asynchronous) and $r(0)\approx 1.0$ (i.e., highly synchronous), respectively, with the results demonstrated in Fig.
-4 for $N=50$.
+4(h) and 4(i)), whereas, for $\epsilon_{2}=6.0$, the system does not undergo any de-synchronisation transition upon reducing $E$; see Fig.
+4(j).
+3.3 Effect of finite system size
+Real bio-physical systems, such as the brain, are typically modelled as a connected system with a finite number of constituent dynamical units (here, oscillators), the size of which can vary by different orders depending on the level of abstraction.
+Thus, while the previous sections demonstrate steady-state results for an infinitely large number of oscillators, the synchronisation phenomena in finite-sized systems are typically affected by fluctuations which can be better characterised by constructing a probability density of the order parameter, denoted by $p(r)$.
+This is computed by simulating the system for a large number of time-steps (here, $10^{5}$) for a fixed set of system parameters.
+This is subsequently repeated over $100$ initial conditions, such that the first $50$ initial conditions correspond to $r(0)\approx 0.0$ (blue curve), and the remaining correspond to $r(0)\approx 1.0$ (red curve), with the results for $N=50$ shown in Fig.
+5.
 (a) $\epsilon_{1}=0.6$
@@
 (d) $\epsilon_{1}=2.6$
-Figure 4: (a)-(d): Distribution of order parameter $p(r)$, collapsed across time, for $N=50$, and different values of $\epsilon_{1}$.
+Figure 5: (a)-(d): Numerically constructed distribution of the order parameter, i.e., $p(r)$, collapsed across time, for $N=50$, and different values of $\epsilon_{1}$.
 The blue and red curves denote completely asynchronous and synchronous initial conditions, respectively, whereas the vertical line denotes $r_{s}$.
@@
 It is observed that, for $\epsilon_{1}=0.6$ in Fig.
-4(a), both distributions show a bimodal structure, with peaks near $r\approx 0.1$ and $r\approx 0.8$ respectively.
+5(a), both distributions show a bimodal structure, with peaks near $r\approx 0.1$ and $r\approx 0.8$, respectively.
 However, the blue curve is substantially skewed towards $r=0$, whereas the red curve has two peaks of similar height.
-This indicates that, if initialized asynchronously, there is a higher probability of the system remaining asynchronous compared to transitioning to a highly synchronous state.
-On the other hand, if the system is initialized in a highly synchronous state, there is a nearly equal probability that the system will transition to an asynchronous state or remain in a synchronous state.
+This indicates that, if initialised asynchronously, there is a higher probability of the system remaining asynchronous compared to transitioning to a highly synchronous state.
+On the other hand, if the system is initialised in a highly synchronous state, there is a nearly equal probability that the system will transition to an asynchronous state or remain in a highly synchronous state.
 In comparison, for $\epsilon_{1}=1.3$ in Fig.
-4(b), both distributions retain their bimodal structure.
+5(b), both distributions retain their bimodal structure.
 However, in contrast to the previous case, both distributions are skewed heavily towards $r=1$.
@@
 This implies that, for any initial condition, there is a much higher probability of observing the system to be in a highly synchronous state; see Figs.
-4(c) and 4(d).
-Conclusions — This work investigates the emergence of synchrony in a population of Kuramoto oscillators, where the interactions are both adaptive and higher-order in nature, having both rotational symmetry-preserving and symmetry-breaking components.
-To this end, exact analytical results for the synchronization states, their stability, and the corresponding phase transition points are obtained in the thermodynamic limit for any order of interaction.
-Subsequently, these were verified through large-scale simulations, focusing on the effects of triadic interactions, revealing a repertoire of dynamical states, including monostable, bistable, and unstable cases.
-Furthermore, varying the relative coupling strength between pairwise and higher-order interactions shows a qualitative change in the system dynamics, including a scenario where the desynchronization transition is suppressed.
-The effects of finite system size were also examined through numerically constructed order parameter distributions, highlighting how fluctuations in the temporal variation of the order parameter, arising out of the finite size of the system, can lead to transitions to a highly synchronous state that would not otherwise occur in the thermodynamic limit.
+5(c) and 5(d).
+Taken together, the results presented in Fig.
+5 highlight the effects of size-induced fluctuations in driving the system to emergent states not observed in the infinite-size limit.
+Furthermore, additional plots of $p(r)$, obtained by repeating the same procedure for different values of $\epsilon_{1}$ and $N$, are presented in Fig.
+6 of C, which highlight the qualitative variation in the characteristic order parameter dynamics as a function of system size.
+4 Conclusions & Future Directions
+This work presents a very general theory of synchronization in a population of Kuramoto oscillators, where the interactions have both rotational symmetry-preserving and symmetry-breaking components.
+Furthermore, the additional terms are included such that they capture a wide range of coupling mechanisms, introduced through adaptation functions that modify the instantaneous coupling strength as a function of the synchronization order parameter, as well as the effects of frustration incorporated using a constant phase lag.
+Although these factors have been studied individually in the literature, this work introduces a unified mathematical framework to explore their collective interplay.
+To this end, exact analytical results for the synchronisation states, their stability, and the corresponding phase transition points are obtained in the thermodynamic limit for any order of interaction using the popular Ott-Antonsen Ansatz.
+Subsequently, these analytical results were verified through large-scale numerical simulations, focusing on the combined effects of pairwise and triadic interactions and an adaptation function having a quadratic dependence on the order parameter, revealing a repertoire of dynamical states, including monostable, bistable, and unstable synchronous cases.
+Furthermore, varying the relative coupling strength between pairwise and higher-order interactions shows a qualitative change in the system dynamics, highlighting both continuous and explosive phase transitions, as well as a scenario where the desynchronization transition is suppressed.
+Finally, the effects of finite system size were also examined through numerically constructed order parameter distributions.
+The obtained results highlight how fluctuations in the temporal variation of the order parameter, arising out of the finite size of the system, can lead to transitions to a highly synchronous state that would not otherwise occur in the thermodynamic limit.
+Such transitions occur spontaneously, i.e., without any change in system parameters, and are distinct from those observed in a system of nearly-infinite size, where transitions occur only upon tuning system parameters across a critical transition point.
 The results presented provide important insights into the interplay of higher-order interactions and other coupling modalities, allowing for more precise delineation of transition boundaries, yielding a comprehensive “map” of the accessible states.
-A notable application lies in understanding epileptic seizures, where incorporating higher-order interactions could improve surgical strategies for drug-resistant epilepsy [schramm2008temporal, jirsa2023personalised, wang2023delineating] by revealing underlying epileptogenic foci that can remain obscured in purely pairwise models, thereby facilitating precise and targeted interventions.
-Another important example concerns the dissemination of information in social systems, where analyzing higher-order connectivity can uncover underlying “cells”, i.e., individuals or organizations that act as synchronized sources of misinformation, and could explain why misleading content often propagates far more rapidly than verified information [vosoughi2018spread].
-Future work could extend this framework to involve empirical networks that include hyper-graph connectivities (for example, see Ref.
-[ko2022growth]), integrate domain-specific models beyond the generic Kuramoto framework adopted here, and estimate parameters using available datasets to obtain specific insights.
-Acknowledgements — The authors acknowledge the financial support provided via NBRC Core Funds. DB acknowledges discussions with Dr. Proloy Das (NBRC), PV Vinitha (NBRC), and Shashwati Tripathi (NBRC).
-Declaration of competing interest — The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Appendix
-Varying system size — In Fig.
-5, the figures highlight the characteristic change in the order parameter dynamics as the system size $N$ is increased for different values of the pairwise coupling strength $\epsilon_{1}$.
+Such models find widespread application across domains, among which a notable application lies in understanding epileptic seizures, where incorporating higher-order interactions of different orders could improve surgical strategies for drug-resistant epilepsy [schramm2008temporal, jirsa2023personalised, wang2023delineating] by revealing underlying epileptogenic foci that can remain obscured in purely pairwise models, thereby facilitating precise and targeted interventions.
+Another important example concerns the dissemination of information in social systems, where analysing higher-order connectivity can uncover underlying “cells”, i.e., individuals or organisations that act as synchronised sources of misinformation, and could explain why misleading content often propagates far more rapidly than verified information [vosoughi2018spread].
+However, in spite of the theoretical advancements presented in this study, certain key limitations still remain.
+This includes obtaining the exact order of interaction driving the explosive spread of synchronous activity in various physical systems, as well as the explicit functional form of the adaptation functions governing the overall order parameter dynamics, thereby opening up many potential directions of future research.
+Furthermore, this will also include extending this framework to involve empirical hyper-graph connectivities (for example, see Ref.
+[ko2022growth]), integrating domain-specific models beyond the generic Kuramoto model, and estimating the various systems parameters using available datasets to obtain problem-specific insights.
+Author Contributions
+DB: Conceptualization (equal); Formal analysis (equal); Investigation (lead); Methodology (equal); Software (lead); Validation (equal); Visualization (lead); Writing – original draft (equal).
+AB: Conceptualization (equal); Funding acquisition (lead); Investigation (supporting); Methodology (supporting); Project administration (lead); Supervision (lead); Visualization (supporting); Writing – review & editing (lead).
+Acknowledgements
+The authors acknowledge the financial support provided via NBRC Core Funds. DB acknowledges discussions with Dr. Proloy Das (NBRC), PV Vinitha (NBRC), and Shashwati Tripathi (NBRC).
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data Availability Statement
+The data used in this article is available online [karmelic2022emergent, bougou2025mesoscale, dataset].
+Appendix A Expression of Jacobians
+The stability of the steady states, denoted by $(r_{s},\psi_{s})$, of Eqs. (14) and (15) can be determined from the eigenvalues of their Jacobian $J$, the elements $J_{ij}$ of which are given by
+$\begin{split}J_{11}={}&\frac{-2r_{s}^{2}\Delta}{(1-r_{s}^{2})}+r_{s}(1-r_{s}^{2})\left[\sum_{d=1}^{D}\left.\frac{{\rm d}\sigma_{d}(r)}{{\rm d}r}\right|_{r_{s}}\cos(\delta_{d})\right.+\\ &\left.\sum_{d=2}^{D}\left.\frac{{\rm d}\sigma_{d}^{\prime}(r)}{{\rm d}r}\right|_{r_{s}}\cos(\chi_{d,s}^{\prime})\right],\end{split}$ (22)
+$\begin{split}J_{21}={}&\frac{-2r_{s}\omega_{0}}{(1+r_{s}^{2})}-(1+r_{s}^{2})\left[\sum_{d=1}^{D}\left.\frac{{\rm d}\sigma_{d}(r)}{{\rm d}r}\right|_{r_{s}}\sin(\delta_{d})\right.+\\ &\left.\sum_{d=2}^{D}\left.\frac{{\rm d}\sigma_{d}^{\prime}(r)}{{\rm d}r}\right|_{r_{s}}\sin(\chi_{d,s}^{\prime})\right],\end{split}$ (23)
+$J_{12}={}$ $2(h-1)r_{s}(1-r_{s}^{2})\sum_{d=2}^{D}\sigma_{d}^{\prime}(r_{s})\sin(\chi^{\prime}_{d,s}),$ (24)
+and
+$J_{22}={}$ $2(h-1)(1+r_{s}^{2})\sum_{d=2}^{D}\sigma_{d}^{\prime}(r_{s})\cos(\chi_{d,s}^{\prime}).$ (25)
+For the choice of parameter values of the simplified model in Sec. 3, the expressions for the elements of the Jacobian reduce to the form
+$\begin{split}J_{11}={}&\frac{-2r_{s}^{2}\Delta}{(1-r_{s}^{2})}+r_{s}(1-r_{s}^{2})\left[\left.\frac{{\rm d}\sigma_{1}(r)}{{\rm d}r}\right|_{r_{s}}+\right.\\ &\left.\frac{1}{\sigma_{2}(r_{s})}\left(\frac{\Delta}{1-r_{s}^{2}}-\sigma_{1}(r_{s})\right)\left.\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}\right|_{r_{s}}\right],\end{split}$ (26)
+$J_{21}={}$ $\frac{-\omega_{0}}{(1+r_{s}^{2})}\left[2r_{s}+\frac{1+r_{s}^{2}}{\sigma_{2}(r_{s})}\left.\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}\right|_{r_{s}}\right],$ (27)
+$J_{12}={}$ $2r_{s}(1-r_{s}^{2})\left[\frac{\omega_{0}}{1+r_{s}^{2}}\right],$ (28)
+and
+$J_{22}={}$ $2(1+r_{s}^{2})\left[\frac{\Delta}{1-r_{s}^{2}}-\Big(\sigma_{1}(r_{s})+\sigma_{2}(r_{s})\Big)\right],$ (29)
+where
+$\sigma_{1}(r)=\frac{\epsilon_{1}}{2\left(1+\gamma r\right)^{2}},\quad\sigma_{2}(r)=\frac{\epsilon_{2}}{2}(1+\gamma r)^{2}r^{2},$ (30)
+and
+$\frac{{\rm d}\sigma_{1}(r)}{{\rm d}r}=\frac{-\gamma\epsilon_{1}}{(1+\gamma r)^{3}},\quad\frac{{\rm d}\sigma_{2}(r)}{{\rm d}r}=\epsilon_{2}r\left[1+\gamma r(3+2\gamma r)\right],$ (31)
+respectively. Note that the expressions depend solely on $r_{s}$ in this case, thus eliminating the need to separately calculate $\psi_{s}$, thereby simplifying calculations.
+Appendix B Simulation details
+A direct simulation of the Kuramoto model, such as those with repeated initial conditions and different parameter values (i.e., Figs. 4 and 5), with a coupling term of the form
+$I=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}+\sigma\phi_{i}+\alpha\right)$ (32)
+would be prohibitively time-consuming due to the presence of nested summations, especially for $d\geq 2$.
+To address this, the RHS of Eq.
+(32) is expanded as $I=X_{d}\cos(\sigma\phi_{i}+\alpha)+Y_{d}\sin(\sigma\phi_{i}+\alpha)$, where
+$X_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right),\quad Y_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\cos\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right).$ (33)
+Therefore, $X_{d}$ can be expanded as
+$X_{d}=\sum_{j_{1},..,j_{d}=1}^{N}\left[\sin\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\cos(c_{d}\phi_{j_{d}})+\right.\\ \left.\cos\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\sin(c_{d}\phi_{j_{d}})\right],$ (34)
+which implies $X_{d}=X_{d-1}C_{d}+Y_{d-1}S_{d}$ and
+$C_{d}=\sum_{j=1}^{N}\cos(c_{d}\phi_{j}),\quad S_{d}=\sum_{j=1}^{N}\sin(c_{d}\phi_{j}).$ (35)
+Similarly, $Y_{d}=Y_{d-1}C_{d}-X_{d-1}S_{d}$, whereas $(X_{1},Y_{1})=(S_{1},C_{1})$.
+This provides an iterative method to efficiently compute the interaction terms of the form of Eq.
+(32), and can be extended to Kuramoto-like coupling terms in any dynamical systems.
+Unless mentioned otherwise, the simulations of Eq.
+(18) were performed using the popular 4th-order Runge-Kutta algorithm [butcher1996history], for a system of $N=10^{4}$ oscillators for $10^{4}$ time-steps, with each individual time-step being $10^{-2}$ seconds.
+The algorithm was implemented in C++, using OpenMP for parallelisation of the computationally intensive portions [dagum1998openmp].
+The solutions of Eq.
+(21) were estimated using the function fsolve(), whereas eigenvalues of the Jacobians in A were obtained using the function eig() in MATLAB.
+All the computations were executed on a workstation equipped with an Intel Core i9-14900K processor and 64 GB DDR5 RAM.
+Appendix C Varying system size
+In Fig. 6, the figures highlight the characteristic change in the order parameter dynamics as the system size $N$ is increased for different values of the pairwise coupling strength $\epsilon_{1}$.
 Figure 6: (a)-(l): $p(r)$ for different values of $(N,\epsilon_{1})$, where $N=100$, $500$, $1000$ (left to right), and $\epsilon_{1}=0.6$, $1.3$, $1.8$, $2.6$ (top to bottom).
 Here, $(\omega_{0},\epsilon_{2},\gamma)=(1.5,5.0,0.2)$, and the colours carry the same meaning as Fig.
-4.
+5.
 The results demonstrate that, for values of $\epsilon_{1}$ given by $0.6$, $1.3$, and $1.8$, the blue curve (asynchronous initial conditions) loses its bi-modal structure and becomes sharply peaked near $r=0$ as the value of $N$ is increased; see Fig.
-5(a)-5(i).
+6(a)-6(i).
 In contrast, for $\epsilon_{1}=2.6$, the blue curve becomes sharply peaked near $r\approx 0.8$ upon increasing $N$; see Figs.
-5(j)-5(l).
+6(j)-6(l).
 These observations are consistent with the fact that, in the limit $N\rightarrow\infty$, the system has a stable fixed point at $r=0$ $\forall$ $\epsilon_{1}\leq 2$.
@@
 Finally, upon increasing $\epsilon_{1}$ to $1.8$ and $2.6$, the red curve becomes unimodal with a sharp peak near $r\approx 0.8$ as the value of $N$ is increased, which is indicative of a stable synchronous fixed point.
-Simulation details — A direct simulation of the Kuramoto model with a coupling term of the form
-$I=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}+\sigma\phi_{i}+\alpha\right)$ (11)
-would be prohibitively time-consuming due to the presence of nested summations, especially for $d\geq 2$.
-To address this, the RHS of Eq.
-(11) is expanded as $I=X_{d}\cos(\sigma\phi_{i}+\alpha)+Y_{d}\sin(\sigma\phi_{i}+\alpha)$, where
-$X_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\sin\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right),\ Y_{d}=\sum_{\mathclap{j_{1},..,j_{d}=1}}^{N}\cos\left(\sum_{k=1}^{d}c_{k}\phi_{j_{k}}\right).$ (12)
-Therefore, $X_{d}$ can be expanded as
-$X_{d}=\sum_{j_{1},..,j_{d}=1}^{N}\left[\sin\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\cos(c_{d}\phi_{j_{d}})+\right.\\ \left.\cos\left(\sum_{k=1}^{d-1}c_{k}\phi_{j_{k}}\right)\sin(c_{d}\phi_{j_{d}})\right],$ (13)
-which implies $X_{d}=X_{d-1}C_{d}+Y_{d-1}S_{d}$ and
-$C_{d}=\sum_{j=1}^{N}\cos(c_{d}\phi_{j}),\ S_{d}=\sum_{j=1}^{N}\sin(c_{d}\phi_{j}).$ (14)
-Similarly, $Y_{d}=Y_{d-1}C_{d}-X_{d-1}S_{d}$, whereas $(X_{1},Y_{1})=(S_{1},C_{1})$. This provides an iterative method to efficiently compute the interaction terms of the form of Eq. (11).